I need help. It's been a while and I need to be refreshed. I know the answer is 5x + 3 .... but I don't know how to get there!
Problem: (x^2+3x-2)-(x^2-2x-5)=

Answer by tutorcecilia(2152)

x^2+3x-2
-(x^2-2x-5)
____________
0x^+5x+3
or 5x+3 because:
3x-(-2x)=3x+2x=5x
and -2-(-5)=-2+5=3